# OM4OV: Leveraging Ontology Matching for Ontology Versioning

### v2

@@ III-B Measurement
OV can reuse these measures, but they need to be extended into four sub-measures for add, delete, remain, and update performance. Within each sub-measure, the precision-recall trade-off still holds. Across different sub-measures, they are not independent but satisfy the following equations, where $N(O)$ is the number of entities in $O$ and $N(O^{\prime})$ is the number of entities in $O^{\prime}$:
@@
$|\Delta A|=|\Delta A\odot|+|\Delta A\otimes|=-|\Delta A\oplus|=-|\Delta A\ominus|$ (7)
We define the corresponding changes of $|A\cap R|$ in remain, update, add, and delete as $\Delta(A\cap R)\odot\otimes$ and $\Delta(A\cap R)\oplus\ominus$.
  • For recall, there are direct relations between different sub-measures. Recalldirect-$\odot$ and Recalltensor-$\otimes$ will increase with Recalldirect-$\oplus$ and Recallsymmetric-$\ominus$ decreasing and vice versa.
    $Recall\odot\otimes\uparrow=\frac{|A\cap R|+|\Delta(A\cap R)\odot\otimes|\uparrow}{|R|}$ (8)
    $Recall\oplus\ominus\downarrow=\frac{|A\cap R|-|\Delta(A\cap R)\oplus\ominus|\downarrow}{|R|}$
@@
  • For precision, there are no direct relations between different sub-measures. The indirect relations, described qualitatively in the following as nondecreasing ($\uparrow$), nonincreasing ($\downarrow$), or either (?)?(?)( ? ), will depend on $\frac{|A\cap R|\times\Delta A}{\Delta(A\cap R)\times|A|}$ in each sub-measure.
    $Precision\odot\otimes\ ?=\frac{|A\cap R|+|\Delta(A\cap R)\odot\otimes|\uparrow}{|A|+|\Delta A|\uparrow}$ (9)
    $\frac{|A\cap R|\times|\Delta A|}{|\Delta(A\cap R)\odot\otimes|\times|A|}<1\quad Precision\odot\otimes\uparrow$
    $\frac{|A\cap R|\times|\Delta A|}{|\Delta(A\cap R)\odot\otimes|\times|A|}>1\quad Precision\odot\otimes\downarrow$
    $Precision\oplus\ominus\ ?=\frac{|A\cap R|-|\Delta(A\cap R)\oplus\ominus|\downarrow}{|A|-|\Delta A|\downarrow}$
    $\frac{|A\cap R|\times|\Delta A|}{|\Delta(A\cap R)\oplus\ominus|\times|A|}>1\quad Precision\oplus\ominus\uparrow$
    $\frac{|A\cap R|\times|\Delta A|}{|\Delta(A\cap R)\oplus\ominus|\times|A|}<1\quad Precision\oplus\ominus\downarrow$
III-C Testbed
There is a dearth of lab-based datasets for OV evaluation. The Ontology Alignment Evaluation Initiative (OAEI) contains several datasets related to OM tasks, but none of them contain benchmarks designed for OV tasks. We propose an approach to constructing synthetic OV datasets from OM datasets. Fig. 2 illustrates the generation of OM4OV datasets. Our approach is described in the following steps:
Figure 2: Generation of OM4OV datasets.
  1. 1.
  2. 2.
    There are four possible changes in OV tasks: remain, update, add, and delete. Each ontology entity in $O_{i}$ is randomly assigned one of these. For notational convenience henceforth, we will treat each element of remain, add, or delete to be either a single entity $e$ or equivalently the idempotent mapping $(e,e)$. Elements of update are necessarily mappings $(e,e^{\prime})$ (also written $e\rightarrow e^{\prime})$ where $e\neq e^{\prime}$ and $e^{\prime}$ is the updated entity name of $e$, but when we write $e\in\textit{update}$ we mean $(e,e^{\prime})\in\textit{update}$ for some $e^{\prime}$. By construction, we also have the four sets as pairwise disjoints.
  3. 3.
    For entities assigned to update, we need to generate the updated entity name. We should expect the new name to have a similar meaning to its original name. For example, the entity name “ConferenceVenuePlace” could be replaced with “Conference_hall” or “Conference_building”, but not the general names “Place” or “Location”. To achieve this goal, we retrieve all equivalent entities provided by reference.xml included in the original OAEI datasets and use them as a replacement synonym corpus. For those entities whose names are unique identifiers or codes (and not textually-meaningful names), we use their annotation properties (e.g. rdfs:label, rdfs:comment, skos:prefLabel, and skos:definition) instead. For those entities that do not have synonyms in the generated list, we randomly re-assign the entity to remain, add, or delete.
  4. 4.
  5. 5.
    We generate $O$ and $O^{\prime}$ according to the following two rules:
    1. (a)
    2. (b)
  6. 6.
    We also incorporate the cross-reference mechanism into our OV testbed, $R_{or}$ and $R_{o^{\prime}r}$ are created according to the following rules: (1) $R_{or}$ is the original reference.xml that removes all the mappings related to add entities. (2) $R_{o^{\prime}r}$ is the original reference.xml that removes all the mappings related to delete entities and updates all the mappings related to update entities.
Unlike the original OAEI datasets for OM, randomness ensures that the synthetic OAEI datasets for OV are different each time they are constructed. This suits the dynamic nature of OV, where the changes vary between different versions. For this reason, we consider the new OAEI datasets for OV more like a testbed, as they can simulate a variety of situations for OV tasks. The current version of the OV testbed contains a total of 12 distinct ontologies from three different OAEI tracks. Table I lists the detailed information of the selected OAEI track used for the OV testbed. The anatomy track contains two large ontologies, while the MSE track has three medium ontologies, and the conference track has 7 small ontologies.
TABLE I: Selected OAEI track for the OV testbed.
Track Domain Number of Ontologies Anatomy Human and Mouse Anatomy 2 Conference Research Conference 7 MSE Materials Science & Engineering 3
III-D Cross-Reference Mechanism
@@
Often, ontology creators provide cross-references to other ontologies to enhance interoperability. For example, the cross-reference between the CMT ontology and the Conference ontology is provided along with the CMT ontology. Reusing these cross-references developed for OM tasks, we propose a novel mechanism to reduce matching candidates and also to improve overall OV performance.
@@
Fig. 3 illustrates the cross-reference mechanism used in the OM4OV pipeline. We can see that, without using the cross-reference $O_{r}$, the matching candidates cover the range of $O\cup O^{\prime}$. This number can be significantly reduced by removing prior matches (i.e. $O\cap O_{r}\cap O^{\prime}$) and non-matches (i.e. $O\cap O_{r}\setminus O^{\prime}$ and $O^{\prime}\cap O_{r}\setminus O$). The prior matching will be part of the final alignment, while the known non-matching will be completely removed in the subsequent OV process. The OV process then only determines the posterior alignment. In practice, the prior alignment usually contains a large number of remain entities and a small number of update entities. Matching performance is also improved by using these known mappings. Since prior matches are inferred from the OM references validated by domain experts, they represent a solid ground truth for alignment in a specific domain. On the other hand, while the known non-matches are removed from the OV process, this reduces the complexity of detecting the posterior alignment.
@@
$R_{or}$ =$=\{(e1,e3,\equiv,c)|e1,e3\in O_{r}\cap O,c\geq s\}$ (10)
$R_{o^{\prime}r}$ =$=\{(e2,e3,\equiv,c)|e2,e3\in O_{r}\cap O^{\prime},c\geq s\}$
We can use $R_{or}$ and $R_{o^{\prime}r}$ to infer some known mappings between $O$ and $O^{\prime}$ before performing OV. We call these mappings a prior alignment ($A_{\pi}$). After subsequently performing OV, we have our posterior alignment ($A_{\pi^{*}}$). Therefore, $A_{match}$ in OV can be decomposed into two parts:
@@
  1. 1.
    $A_{\pi}$ can be directly inferred from the two cross-references $R_{or}$ and $R_{o^{\prime}r}$. The equivalence relation is transitive, so for $e1\in O$, $e2\in O^{\prime}$, and $e3\in O_{r}$, if $e1=e3$ and $e2=e3$ then $e1=e2$.
    $A_{\pi}=R_{or}\cap R_{o^{\prime}r}=\{(e1,e2,\equiv,c)|e1,e2\in O\cap O_{r}\cap O^{\prime},c\geq s\}$ (12)
  2. 2.
    $A_{\pi^{*}}$ aims to detect missing mappings from the cross-reference. None of these mappings would come from any pairwise intersection of $O$, $O_{r}$, and $O^{\prime}$ because $O\cap O_{r}\setminus O^{\prime}$ and $O^{\prime}\cap O_{r}\setminus O$ are pre-defined as non-matched entities, and the matched entities in $O\cap O_{r}\cap O^{\prime}$ have already been captured in the $A(\pi)$. As a result, $A_{\pi^{*}}$ can be defined within a smaller scope:
    $A_{\pi^{*}}=\{(e1,e2,\equiv,c)|e1\in O\setminus O_{r},e2\in O^{\prime}\setminus O_{r},c\geq s\}$ (13)
@@
$A_{\pi^{*}}=\{(e1,e2,\equiv,c)|e1\in O\setminus O_{ra},e2\in O^{\prime}\setminus O_{ra},c\geq s\}$
$\text{where}\ O_{ra}=O_{r1}\cup O_{r2}\cup...O_{rn}$
IV Implementation & Evaluation
The OM4OV pipeline is implemented in Agent-OV, a variant of Agent-OM [14]. Agent-OM is an agent-powered LLM-based OM system. Its foundation framework is designed for traditional OM tasks. We extend the original framework with the OM4OV pipeline so that it can be used to handle OV tasks.
As an extension of Agent-OM, Agent-OV also supports a wide range of LLMs, including commercial API-accessed LLMs OpenAI GPT [15], Anthropic Claude [16], and Mistral AI [17], as well as open-source LLMs Meta Llama [18], Google Gemma [19], and Alibaba Qwen [20]. For the performance of Agent-OV using different LLMs, we refer the reader to [14], where we find that API-accessed LLMs generally perform better than open-source LLMs.
IV-A Evaluation of OV testbed
Fig. 4 shows the evaluation of the Agent-OV system on the OV testbed. We use the OpenAI model gpt-4o-mini (2024-07-18) [21] for evaluation. The embedding chosen correspondingly is OpenAI text-embedding-ada-002 [22]. The hyperparameter settings are $similarity\_threshold=0.90$ and $top@k=3$ across all alignments generated. We run the OV testbed for 5 iterations, simulating 5 distinct versions for each ontology. The results indicate that it is possible to unify the OM and OV tasks. With the necessary modifications, the OM systems can also be used in OV tasks.
(a) Anatomy Track: human
(b) Anatomy Track: mouse
(c) Conference Track: cmt
(d) Conference Track: conference
(e) Conference Track: confof
(f) Conference Track: edas
(g) Conference Track: ekaw
(h) Conference Track: iasted
(i) Conference Track: sigkdd
(j) MSE Track: EMMO
(k) MSE Track: MatInfo
(l) MSE Track: MatOnto
Figure 4: Evaluation of OV testbed.
  1. 1.
    Our experiments show that the OM system achieves acceptable performance in tracking changes over different versions of the ontologies. Interestingly, the performance of our system on OV tasks is generally better than on OM tasks. This could be due to the fact that the ontologies used in the OV tasks are two versions of the same ontology that share consistent ontology design patterns.
  2. 2.
    For the four micro-level sub-measures, we observe:
    1. (a)
      The performance is highest in remain, followed by add and delete, and it is relatively low in update. This trend is consistent across different tracks and ontologies.
    2. (b)
      The measurements for remain are commonly very close to 100% and not statistically significant in OV tasks.
    3. (c)
      The measures for add and delete are generally good. This may be because our system uses LLMs as the backend, and LLMs generally have strong background knowledge to detect missing mappings.
    4. (d)
      The measurements for update show scope for improvement. This is because finding non-trivial alignments and appropriate similarity thresholds is not easy. Unlike for OM tasks, we believe mixing remain and update in OV tasks is a common pitfall and can cause skewed performance measurement.
  3. 3.
    For the computational time, we observe a longer computation time for the large-scale ontologies in the Anatomy Track. Although Agent-OV has an optimisation module for the matching candidate selection process (inherited from Agent-OM), it is still not sufficient for some OV tasks.
IV-B Evaluation of cross-reference mechanism
@@
Fig. 5 shows the comparison of the OV performance with and without using the cross-reference (CR) mechanism on the same OV testbed we analysed in Section IV-A. We can see that the CR mechanism significantly improves the capability of Agent-OV to detect update entities and slightly improves the performance on detecting add and delete entities, with no effect on the remain entities. More specifically, the performance improvement is mainly caused by the recall for update entities, while this is more evident in the precision for add and delete entities.
(a) Remain-Precision
(b) Update-Precision
(c) Add-Precision
(d) Delete-Precision
(e) Remain-Recall
(f) Update-Recall
(g) Add-Recall
(h) Delete-Recall
(i) Remain-F1 Score
(j) Update-F1 Score
(k) Add-F1 Score
(l) Delete-F1 Score
@@
Figure 5: Evaluation of cross-reference (CR) mechanism.
IV-C Evaluation of hyperparameter settings
Fig. 6 shows the comparison of using different similarity thresholds in the CMT Ontology versioning. Similar to OM tasks, changes in hyperparameter settings in OV tasks will lead to a trade-off in precision and recall. Moreover, the hyperparameter settings also influence the sub-measures. We find that the similarity threshold has no effect on detecting remain entities, but significantly affects the detection of update entities and slightly affects the detection of add and delete entities. Also, detecting update entities and detecting add and delete is negatively correlated. Lower similarity thresholds can result in more update entities being detected, while higher similarity thresholds may find more add and delete entities. The result of the experiment also shows that our cross-reference mechanism is less sensitive to the similarity threshold, which means that it can be useful in scenarios where the optimal similarity threshold is unclear or hard to determine.
Figure 6: Evaluation of different similarity thresholds.
V Discussion
@@
  1. 1.
    False mappings in OV can be a wrong ontology design choice. In the following example, the reference shows that cmt:writtenBy is updated to cmt:isWrittenBy, but the OV system will consider this entity to remain unchanged because a false mapping (cmt:writtenBy, cmt:writtenBy, $\equiv$, 1) is detected in the remain subset. This issue is caused by incorrect name conventions. In the CMT ontology, cmt:hasAuthor and cmt:writtenBy are different entities because one is targeting the paper (cmt:Paper, cmt:hasAuthor, cmt:Author) and one is targeting the review (cmt:review, cmt:writtenBy, cmt:reviewer). However, using cmt:writtenBy for paper still makes sense (cmt:Paper, cmt:writtenBy, cmt:Author). Within one ontology, the meaning of two entities is too close to be distinguished and therefore leads to them being synonyms for each other. Ideally, we should avoid this type of ontology design. This example also demonstrates a unique benefit of using OM for OV, which could potentially assist in ontology design.
@@
  2. 2.
    False mappings in OV can be an ambiguous “equivalent” relationship. In the following example, the reference shows that cmt:ConferenceChair can be updated to cmt:General_Chair to indicate a specific type of chair responsible for coordinating the conference. However, the OV system may predict that cmt:ConferenceChair is equivalent to cmt:Chair. This interpretation is not wrong but follows a different ontology design pattern. It is important to notice that the term ”equivalent” in OV is weaker than that in OM. OV allows for roughly “equivalent” mappings. The entities mapped in OV can slightly alter their meanings in response to changes in the domain knowledge understanding.
@@
  3. 3.
    False mappings in OV can be an inappropriate setup on the similarity threshold. In the following example, if $similarity\_threshold\ (s)=0.95$, cmt:SubjectArea in $O$ will be assigned to delete entities as it does have matching entities; If $similarity\_threshold\ (s)=0.90$, cmt:SubjectArea in $O$ will be assigned to update entities as it has a matching entity cmt:Topic in $O^{\prime}$. Both results are valid because defining the boundary between matching and non-matching is context- and application-dependent. For example, the similarity threshold is relatively higher in the biomedical domain to ensure each terminology is unique, whereas the similarity threshold in the conference domain can be relatively lower to improve the interoperability of terminologies used in research conferences.
@@
VI Limitations
We focus mainly on tracking conceptual changes in classes and properties (e.g. adding, deleting, or updating). In practice, there are also internal relationship changes (e.g. changing the domain and range of a class or moving a sibling class to a different parent). These changes are currently only indicated by the changes in the similarity score, while details can only be observed by inspecting the classes and properties. Our future work aims to improve the explainability of changes.

### v3

@@ 4.3 Experimental Results
Within each sub-measure, the precision-recall trade-off still holds. Across different sub-measures, they are not independent but satisfy the following equations, where $N(O)$ is the number of entities in $O$ and $N(O^{\prime})$ is the number of entities in $O^{\prime}$:
@@
$|\Delta A|=|\Delta A\odot|+|\Delta A\otimes|=|\Delta A\oplus|=|\Delta A\ominus|$ (6)
@@
We define the corresponding changes of $|A\cap R|$ in remain, update, add, and delete as $\Delta(A\cap R)\odot\otimes$ and $\Delta(A\cap R)\oplus\ominus$. For recall, there are direct relations between sub-measures. Recalldirect-$\odot$ and Recalltensor-$\otimes$ will increase with Recalldirect-$\oplus$ and Recallsymmetric-$\ominus$ decreasing and vice versa. For precision, there are no direct relations between sub-measures. The indirect relations, described qualitatively in the following as increasing ($\uparrow$), decreasing ($\downarrow$), or either (?)?(?)( ? ), will depend on $\frac{|A\cap R|\times\Delta A}{\Delta(A\cap R)\times|A|}$ in each sub-measure.
$Rec\odot\otimes\uparrow=\frac{|A\cap R|+|\Delta(A\cap R)\odot\otimes|\uparrow}{|R|}\quad Prec\odot\otimes\ ?=\frac{|A\cap R|+|\Delta(A\cap R)\odot\otimes|\uparrow}{|A|+|\Delta A|\uparrow}$ (7)
$Rec\oplus\ominus\downarrow=\frac{|A\cap R|-|\Delta(A\cap R)\oplus\ominus|\downarrow}{|R|}\quad Prec\oplus\ominus\ ?=\frac{|A\cap R|-|\Delta(A\cap R)\oplus\ominus|\downarrow}{|A|-|\Delta A|\downarrow}$
5 Pipeline Optimisation
5.1 Motivation
@@
Often, ontology creators provide cross-references to other ontologies to enhance interoperability and facilitate integration. For example, a cross-reference between the CMT ontology and the Conference ontology is provided alongside the CMT ontology. Reusing the cross-references developed for OM tasks, we propose a novel mechanism to reduce the number of matching candidates and also to improve overall OV performance.
5.2 Approach
@@
Figure 6 illustrates the cross-reference mechanism used in the OM4OV pipeline. We can see that, without using the cross-reference $O_{r}$, the matching candidates cover the range of $O\cup O^{\prime}$. This number can be significantly reduced by removing prior matches (i.e. $O\cap O_{r}\cap O^{\prime}$) and non-matches (i.e. $O\cap O_{r}\setminus O^{\prime}$ and $O^{\prime}\cap O_{r}\setminus O$). The prior matching will be incorporated into the final alignment, while the known non-matching regions will be removed entirely in the subsequent OV process. The OV process then only determines the posterior alignment. In practice, the prior alignment usually contains a large number of remain entities and a small number of update entities. Matching performance is also enhanced by utilising these established mappings. Since prior matches are inferred from the OM references validated by domain experts, they represent a solid ground truth for alignment in a specific domain. On the other hand, while the known non-matches are removed from the OV process, this simplifies the detection of the posterior alignment.
@@
$R_{or}$ =$=\{(e1,e3,\equiv,c)|e1,e3\in O_{r}\cap O,s\leq c\leq 1\}$ (8)
$R_{o^{\prime}r}$ =$=\{(e2,e3,\equiv,c)|e2,e3\in O_{r}\cap O^{\prime},s\leq c\leq 1\}$
Definition 5.2.
We can use $R_{or}$ and $R_{o^{\prime}r}$ to infer some known mappings between $O$ and $O^{\prime}$ before performing OV. We call these mappings a prior alignment ($A_{\pi}$). After subsequently performing OV, we have our posterior alignment ($A_{\pi^{*}}$). In this setting, $A_{match}$ in OV can be decomposed into two parts: $A_{\pi}$ and $A_{\pi^{*}}$.
@@
Definition 5.2.1.
$A_{\pi}$ can be directly inferred from the two cross-references $R_{or}$ and $R_{o^{\prime}r}$. The equivalence relation is transitive, so for $e1\in O$, $e2\in O^{\prime}$, and $e3\in O_{r}$, if $e1=e3$ and $e2=e3$ then $e1=e2$. $A_{\pi^{*}}$ aims to detect missing mappings from the cross-reference. None of these mappings would come from any pairwise intersection of $O$, $O_{r}$, and $O^{\prime}$ because $O\cap O_{r}\setminus O^{\prime}$ and $O^{\prime}\cap O_{r}\setminus O$ are pre-defined as non-matched entities, and the matched entities in $O\cap O_{r}\cap O^{\prime}$ have already been captured in the $A(\pi)$. As a result, $A_{\pi^{*}}$ can be defined within a smaller scope:
$A_{\pi}$ $=R_{or}\cap R_{o^{\prime}r}=\{(e1,e2,\equiv,c)|e1,e2\in O\cap O_{r}\cap O^{\prime},s\leq c\leq 1\}$ (10)
$A_{\pi^{*}}$ =$=\{(e1,e2,\equiv,c)|e1\in O\setminus O_{r},e2\in O^{\prime}\setminus O_{r},s\leq c\leq 1\}$
Definition 5.2.2.
@@
$A_{\pi^{*}}$ =$=\{(e1,e2,\equiv,c)|e1\in O\setminus O_{ra},e2\in O^{\prime}\setminus O_{ra},c\geq s\}\ \text{where}\ O_{ra}=O_{r1}\cup O_{r2}\cup...O_{rn}$
5.3 Evaluation
@@
Figure 7 shows the comparison of the OV performance with and without the cross-reference (CR) mechanism on the same OV testbed that we analysed in Section 4. We can see that the CR mechanism significantly enhances the capability of Agent-OV to detect update entities and slightly improves performance in detecting add and delete entities, with no effect on remain entities. More specifically, performance improvement is mainly attributed to recall for update entities, while this is more evident in the precision for add and delete entities.
Figure 7: Evaluation of the cross-reference (CR) mechanism on the OV testbed.
Figure 8 shows the comparison of using different similarity thresholds in the CMT Ontology versioning. The results of the experiment also show that our cross-reference mechanism is less sensitive to the similarity threshold, which means it can be helpful in scenarios where the optimal similarity threshold is unclear or difficult to determine.
@@
Figure 8: Evaluation of the cross-reference (CR) mechanism on different similarity thresholds.
6 Discussion
@@
{bracketenumerate}
False mappings in OV can result from a flawed ontology design choice. In the following example, the reference shows that cmt:writtenBy is updated to cmt:isWrittenBy, but the OV system will consider this entity to remain unchanged because a false mapping (cmt:writtenBy, cmt:writtenBy, $\equiv$, 1) is detected in the remain subset. This issue is caused by incorrect name conventions. In the CMT ontology, cmt:hasAuthor and cmt:writtenBy are different entities because one is targeting the paper (cmt:Paper, cmt:hasAuthor, cmt:Author) and one is targeting the review (cmt:review, cmt:writtenBy, cmt:reviewer). However, using cmt:writtenBy for paper still makes sense (cmt:Paper, cmt:writtenBy, cmt:Author). Within one ontology, the meaning of two entities is too close to be distinguished and therefore leads to them being synonyms for each other. Ideally, we should avoid this type of ontology design. This example also demonstrates a unique benefit of using OM for OV, which could potentially assist in ontology design.
@@
False mappings in OV can create an ambiguous “equivalent” relationship. In the following example, the reference shows that cmt:ConferenceChair can be updated to cmt:General_Chair to indicate a specific type of chair responsible for coordinating the conference. However, the OV system may predict that cmt:ConferenceChair is equivalent to cmt:Chair. This interpretation is not wrong, but follows a different ontology design pattern. It is vital to notice that the term “equivalent” in OV is weaker than that in OM. OV allows for roughly “equivalent” mappings. The entities mapped in OV can slightly alter their meanings in response to changes in the domain knowledge understanding.
@@
False mappings in OV can be an inappropriate setting of the similarity threshold. In the following example, if $similarity\_threshold\ (s)=0.95$, cmt:SubjectArea in $O$ will be assigned to delete entities as it does have matching entities; If $similarity\_threshold\ (s)=0.90$, cmt:SubjectArea in $O$ will be assigned to update entities as it has a matching entity cmt:Topic in $O^{\prime}$. Both results are valid because defining the boundary between matching and non-matching is context- and application-dependent. For example, the similarity threshold is relatively higher in the biomedical domain to ensure each terminology is unique, whereas the similarity threshold in the conference domain can be relatively lower to improve the interoperability of terminologies used in research conferences.
@@
7 Limitations
{bracketenumerate}
We focus mainly on tracking conceptual changes in classes and properties (e.g. adding, deleting, or updating class and properties). In practice, there are also internal relationship changes (e.g. changing the domain and range of a class or moving a sibling class to a different parent). These changes are currently only indicated by the changes in the similarity score, while details can only be observed by inspecting the classes and properties. Our future work aims to improve the explainability of changes.